• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 07:52:36 on January23,2009

    Entry number 257405

    keyword=Owl shift summary

    Shift started with no beam in the Hall...

    1:00 MCC called. Want to send 2uA CW beam into the Hall.
    1:25 MCC ready to send beam.
    1:30 BPM reading off by ~1mm, asked MCC to adjust. Sending some pulsed beam.
    1:45 Beam position ok, moving to He3 target
    1:50 Jin came and checked raster size. It was larger than usual (see corresponding entries). He asked for a harp scan to check beam size.

    Beam size ok, and raster ok too after the harp scan, so we resumed data taking at around 5:30.

    Smooth running from then on...